# MSACL: Multi-Step Actor-Critic Learning with Lyapunov Certificates for Exponentially Stabilizing Control

### v1

@@ IV-C Automated Temperature Adjustment
where $\bar{\mathcal{H}}$ denotes the target entropy, typically set as $\bar{\mathcal{H}}=-m$, with $m$ representing the dimension of the action space.
By integrating the loss functions for the soft Q-network $Q_{\theta}$, the Lyapunov network $V_{\psi}$, the policy network $\pi_{\phi}$, and the temperature coefficient $\alpha$, we iteratively update the parameters within an actor-critic framework using the Adam optimizer [37]. The complete algorithmic procedure is summarized in Algorithm 1, and the overall flowchart is illustrated in Fig. 2.
Figure 2: Flowchart of the proposed MSACL algorithm.
@@
   Update $Q_{\theta_{i}}$ via $\theta_{i}\leftarrow\theta_{i}-\beta_{\theta}\nabla_{\theta_{i}}\mathcal{L}_{\text{SoftQ}}(\theta_{i})$ 11footnotemark: 1
   if $\left(\text{iteration count}\mod d\right)=0$ then 22footnotemark: 2
@@
In the flowchart of Algorithm 1, we incorporate two implementation refinements. Specifically, regarding footnote 1: for the soft Q-loss function (14), we replace the target value $Q_{\bar{\theta}}$ with $\min\{Q_{\bar{\theta}_{1}},Q_{\bar{\theta}_{2}}\}$ to mitigate overestimation bias, as noted in Remark 1. Regarding footnote 2: we implement delayed updates for both the policy network $\pi_{\phi}$ and the temperature coefficient $\alpha$. However, within each delayed update cycle, we perform $d$ consecutive updates for these parameters as a “delayed update compensation” mechanism to ensure adequate policy improvement. These techniques are adapted from the cleanrl framework [38], and our experiments demonstrate that integrating these refinements significantly enhances overall performance and training stability.
@@ V Experiments
In this section, we conduct a comprehensive experimental evaluation of the MSACL algorithm across six benchmark environments, encompassing both stabilization and high-dimensional nonlinear tracking tasks. We compare our approach against baseline model-free RL algorithms and Lyapunov-based RL algorithms using rigorous performance and stability metrics. The analysis further includes the visualization of learned Lyapunov certificates, extensive robustness and generalization testing under parametric uncertainties and stochastic noise, and a sensitivity study investigating the impact of the multi-step horizon $n$ on the learning process.
V-A Benchmark Environments
Figure 3: Benchmark environments. (a) Uncontrolled VanderPol (Phase Portrait), (b) Pendulum, (c) Planar DuctedFan, (d) Two-link planar robot, (e) SingleCarTracking, (f) QuadrotorTracking.
In this subsection, we present the benchmark environments utilized to evaluate the experimental performance of the proposed and baseline algorithms, encompassing both stabilization and tracking tasks. We categorize these environments by their primary control objectives. For stabilization tasks, the environments include: Controlled VanderPol (hereafter referred to as VanderPol), Pendulum, Planar DuctedFan (DuctedFan), and Two-link planar robot (Two-link). For tracking tasks, the benchmark comprises SingleCarTracking and QuadrotorTracking. As illustrated in Figure 3, comprehensive specifications and dynamical parameters for all environments are provided in Appendix VI-B.
In stabilization tasks, the control objective is to drive the system state to the equilibrium point, i.e., $\mathbf{x}_{g}=\mathbf{0}\in\mathbb{R}^{n}$. In tracking tasks, the goal is to ensure the system state follows a time-varying reference trajectory $\mathbf{x}_{t}^{\text{ref}}$, which may represent desired coordinates, velocities, or attitudes. Across all tasks, we employ a unified reward structure:
@@ V-A Benchmark Environments
where the first term is a penalty term inspired by the performance functionals commonly utilized in optimal control [39], with $\mathbf{Q_{r}}$ and $\mathbf{R_{r}}$ being positive definite weight matrices. This term penalizes deviations from the equilibrium and excessive control effort at each time step, effectively minimizing the cumulative cost $\sum_{t=0}^{\infty}\left[\mathbf{x}_{t}^{\top}\mathbf{Q_{r}}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R_{r}}\mathbf{u}_{t}\right]$. For tracking tasks, the state $\mathbf{x}_{t}$ in the penalty term is replaced by the tracking error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$. The second term provides sparse encouragement, defined as:
@@
Specifically, when the state approaches the equilibrium (i.e., its infinity norm falls below a small threshold $\delta$, such as 0.01), the environment provides a constant positive reward $r_{\delta}$ (e.g., +1) to further incentivize convergence or minimize the steady-state tracking error. This reward design is intentionally parsimonious and intuitive. Subsequent experiments demonstrate that under this straightforward scheme, our algorithm successfully achieves the control objectives across all benchmark environments.
These six benchmark environments are constructed following the OpenAI Gym framework [40]. Notably, our implementation of the Pendulum environment is significantly more rigorous than the standard OpenAI Gym version. Specifically, we expand the initial state space to $\pi,\pi]\times[-10,10]$ and the action range to [5,5][-5,5], while introducing boundary-based termination conditions and extending the maximum episode length to 1000 steps. These modifications increase the control difficulty, intentionally designed to highlight the performance disparities between different algorithms. Detailed specifications for this and other environments are provided in Appendix VI-B. During each agent-environment interaction, the environment returns a standardized 5-tuple $(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$, where the state transition from $\mathbf{x}_{t}$ to $\mathbf{x}_{t+1}$ is updated using the standard explicit Euler integration method.
@@ V-B Baseline Algorithms
To evaluate the effectiveness of the proposed MSACL algorithm, we compare it against classic model-free RL algorithms, including SAC [19] and PPO [24], as well as state-of-the-art Lyapunov-based RL algorithms, specifically LAC [28] and POLYC [27]. While SAC and PPO have demonstrated superior performance across various benchmarks, they are evaluated here on stabilization and tracking tasks to provide a performance baseline. LAC and POLYC incorporate Lyapunov-based constraints into standard RL frameworks to provide stability guarantees; these approaches serve as primary inspirations for our method.
In the LAC algorithm, the authors replace the standard reward signal with a quadratic cost function, which is defined across environments as:
@@
where $\mathbf{Q_{c}}$ is a positive definite matrix. Notably, this formulation considers only state-related penalties, whereas our reward definition $r_{t}$ explicitly accounts for both state deviations and control effort. By comparing against these baselines, we aim to demonstrate the advantages of MSACL in terms of stability metrics and convergence speed.
To ensure a fair and consistent comparison, all algorithms are implemented within the General Optimal control Problem Solver (GOPS) framework [41], an open-source PyTorch-based RL library. Within this unified architecture, the actor and critic networks utilize identical structures consisting of multi-layer perceptrons (MLPs) with two hidden layers of 256 units each and ReLU activation functions. All actor networks employ stochastic policies parameterized by diagonal Gaussian distributions. Regarding the critic architectures: LAC utilizes only a Lyapunov network, while POLYC incorporates both a value function and a Lyapunov certificate. For fairness, SAC and LAC also implement clipped double Q-learning and delayed policy updates as discussed in Remark 1. Parameter updates for all algorithms are performed using the Adam optimizer. The core hyperparameter configurations are summarized in Table I, with full specifications available in our code repository, which will be publicly released.
@@ V-C Performance Evaluation and Stability Analysis
In this subsection, we present the training outcomes, evaluate stability metrics under the learned policies, and visualize the obtained Lyapunov certificates. To ensure statistical significance, each algorithm was trained using five random seeds per benchmark. The training progress is illustrated in Fig. 4, utilizing the following performance metrics:
  • AMCR (Average Mean Cumulative Reward): The sample mean of the Mean Cumulative Reward (MCR) computed across multiple independent training trials. The MCR is defined as the total undiscounted reward accumulated throughout an episode, normalized by the episode length to provide a consistent per-step measure of policy performance.
  • AMCC (Average Mean Cumulative Cost): The sample mean of the Mean Cumulative Cost (MCC) computed across multiple independent training trials. The MCC is defined as the total undiscounted cost accumulated throughout an episode, normalized by the episode length to quantify the average per-step magnitude of state deviations and control effort.
@@
Figure 4: Training performance across six benchmark environments. The results show the average MCR and MCC over five independent runs. Solid lines and shaded regions represent the mean and one standard deviation, respectively. The horizontal axis indicates the number of training steps.
The training curves demonstrate that MSACL outperforms or matches all baseline algorithms across all environments. Notably, the AMCR for MSACL remains positive and stable during the late training phase. Given our reward design, a positive MCR indicates that the system state has successfully entered the target $\delta$-neighborhood. Higher reward values further imply faster convergence, as the agent spends a greater proportion of the episode receiving the positive reward $r_{\delta}$.
For subsequent comparative analysis, we select the optimal policy model for each algorithm, defined as the specific model that achieved the highest MCR among the five independent training runs. These optimal policies are evaluated from 100 fixed random initial states across all benchmarks. To rigorously assess stability and convergence, we utilize the following metrics across four target radii ($0.2,0.1,0.05,\text{and }0.01$):
  • Reach Rate (RR): The percentage of trajectories that enter the specified radius.
  • Average Reach Step (ARS): The mean number of steps required to first enter the radius.
  • Average Hold Step (AHS): The mean number of steps the state remains within the radius after the first entry.
@@
TABLE II: Experimental Results of Baseline Algorithms Across Benchmark Environments
@@
  • Note: The values for AMCR and AMCC represent the mean performance of the optimal policy (defined as the policy achieving the highest evaluation MCR) averaged over 100 independent trials, with $\pm$ denoting one standard deviation. Stability metrics (RR, ARS, and AHS) are evaluated across four distinct radii, where “–” indicates that the system failed to enter the specified radius within the episode duration. Bold entries highlight the superior results: for performance metrics, the maximum AMCR and minimum AMCC are bolded; for stability metrics, the highest RR is prioritized for bolding, with the shortest ARS and longest AHS selected in the event of identical RR values.
These results highlight the superiority of MSACL. Across all six environments, MSACL consistently achieves the highest AMCR. In terms of AMCC, MSACL attains the absolute minimum in the Pendulum, SingleTrackCar, and QuadrotorTracking tasks, while remaining highly competitive in others. Furthermore, MSACL achieves the highest RR across all radii. Most significantly, in the high-dimensional nonlinear QuadrotorTracking task, MSACL is the only method to achieve a 100% reach rate at the $0.01$ precision level, whereas baselines fail to meet this stringent requirement. These results underscore the robustness and precision of MSACL in complex control applications.
Finally, we visualize the Lyapunov networks $V_{\psi}$ corresponding to the optimal policies. The learned Lyapunov certificates and their associated contour plots are shown in Fig. 5. For high-dimensional systems, the plots represent the relationship between $V_{\psi}$ and two specific state components while setting all other components to zero.
@@
Since MSACL is built upon the SAC framework, SAC serves as the primary baseline for our ablation study. As shown in Fig. 4 and Table II, the integration of the Lyapunov certificate significantly enhances both training performance and stability metrics compared to the standard SAC algorithm. This validates that embedding Lyapunov-based constraints into maximum entropy RL effectively yields neural policies with superior stability guarantees for deterministic control systems.
Figure 5: Visualization of learned Lyapunov certificates and their corresponding contour plots across various benchmark environments: (a) VanderPol; (b) Pendulum; (c)–(e) DuctedFan (projections in the $(x,y)$-plane with $\theta=0$, $(y,\theta)$-plane with $x=0$, and $(x,\theta)$-plane with $y=0$); (f) Two-link; (g) SingleCarTracking; (h)–(j) QuadrotorTracking (projections in the $(e_{x},e_{y})$-plane with $e_{z}=0$, $(e_{y},e_{z})$-plane with $e_{x}=0$, and $(e_{x},e_{z})$-plane with $e_{y}=0$).
@@ V-D Evaluation of Robustness and Generalization
It is well known that neural policies, characterized by high-dimensional parameter spaces, are susceptible to overfitting within specific training environments. This can lead to degraded robustness against stochastic noise and parametric variations, as well as limited generalization to unseen scenarios. In this subsection, we evaluate the robustness and generalization capabilities of the policies trained via MSACL to demonstrate their applicability in real-world environments with inherent uncertainties.
For stabilization tasks (VanderPol, Pendulum, DuctedFan, and Two-link), we introduce process noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\bm{\sigma}_{\text{env}})$ to the system dynamics, resulting in the perturbed form:
@@
Additionally, we modify key physical parameters of the systems to assess whether the MSACL-trained policies can maintain stability under environmental variations.
For tracking tasks (SingleCarTracking and QuadrotorTracking), we evaluate robustness by injecting dynamical noise and assess generalization by requiring the policies to track reference trajectories $\mathbf{x}_{t}^{\text{ref}}$ that differ significantly from those encountered during training. The specific parameter modifications and noise magnitudes utilized for these evaluations are detailed in Table III.
@@
The robustness testing results for the stabilization tasks are illustrated in Fig. 6. Despite the introduction of parametric perturbations and process noise, all system trajectories originating from the specified initial states successfully converge to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$.
The results for the tracking tasks are presented in Fig. 7 and Fig. 8. It is evident that the trained neural policies effectively mitigate noise interference within a reasonable range. In the SingleCarTracking environment, while the policy was trained on straight-line references, we evaluate its generalization using circular and sinusoidal trajectories. Experimental validation demonstrates excellent tracking performance; as shown in the position error plot in Fig. 7e, the tracking error $e_{\mathrm{xy}}=\sqrt{e_{x}^{2}+e_{y}^{2}}$ remains within 0.1 during the steady-state phase.
For the QuadrotorTracking task, the policy was trained using a horizontal helical trajectory. During generalization testing, we evaluate the policy against two unseen trajectories: a vertical helix and a Lissajous curve, as shown in Fig. 8c and Fig. 8d. Notably, the MSACL-trained policy demonstrates superior generalization, with the tracking error $e_{\mathrm{xyz}}=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ stabilizing around 0.02 in the later stages.
Figure 6: System trajectories under the neural policies in stabilization tasks subjected to environmental parametric uncertainties and process noise: (a) VanderPol with state variables $\mathbf{x}=[x,\dot{x}]^{\top}$; (b) Pendulum with angular position $\theta$ and angular velocity $\dot{\theta}$; (c) DuctedFan with coordinates $x,y$ and orientation $\theta$; (d) Two-link with joint angles $q_{1}$ and $q_{2}$. All trajectories exhibit robust convergence to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$ despite external disturbances.
@@
Based on these results, we conclude that policies trained via the MSACL algorithm exhibit robust performance against moderate uncertainties and favorable generalization to unseen reference trajectories. However, our testing also reveals that the policies’ efficacy diminishes under extreme disturbances or environmental parameters that deviate substantially from the training distribution. In such cases of significant domain shift, the MSACL algorithm can be readily redeployed to retrain the policy, ensuring stable and high-precision control for the specific operational scenario.
V-E Sensitivity Analysis of the Multi-Step Horizon $n$
In this subsection, we investigate the influence of the multi-step horizon parameter $n$ on the performance of the MSACL algorithm across all benchmark environments. Following the evaluation protocol detailed in Section V-C, we utilize both performance metrics (AMCR and AMCC) and stability metrics (RR, ARS, and AHS) to assess the policies trained under varying configurations of $n$. Specifically, we conduct 100 randomized experiments for each benchmark by generating and fixing a set of 100 random initial states. Policies trained with different $n$ values interact with the environment starting from these fixed states to compute the aforementioned evaluation metrics.
The results across all environments are summarized in Table IV. The experimental findings reveal a consistent trend: as $n$ increases, the stabilization and tracking performance of the MSACL-trained policies generally improves. This improvement occurs because a larger $n$ provides a longer-horizon trajectory during training, enabling the policy $\pi$ to more accurately identify the stability direction, namely the descent direction of the Lyapunov certificate $V_{\psi}$. This reduces the training bias and reinforces the stability guarantees for the resulting neural policy.
Empirically, setting $n=20$ yields satisfactory performance across all benchmark environments, suggesting it as a versatile default parameter. However, an excessively large $n$ may not always lead to optimal results. As analyzed in Section III-B, longer temporal horizons introduce greater cumulative variance into the training data, which can distort the optimization landscape and result in suboptimal policies. Therefore, selecting an appropriate $n$ is essential to balancing the bias-variance trade-off. In summary, for the control tasks presented in this article, $n=20$ delivers consistently robust results; for specific applications requiring peak performance, fine-tuning around this value is recommended to achieve the optimal balance between training bias and variance.

### v2

@@ IV-C Unified Actor-Critic Optimization Framework
where $\bar{\mathcal{H}}$ denotes the target entropy.
By integrating the loss functions for the soft Q-network $Q_{\theta}$, Lyapunov network $V_{\psi}$, policy network $\pi_{\phi}$, and temperature coefficient $\alpha$, we iteratively update these parameters within an actor-critic framework. The complete procedure is summarized in Algorithm 1 with flowchart illustrated in Fig. 2.
Figure 2: Flowchart of the MSACL algorithm.
@@
   Update $Q_{\theta_{i}}$ via $\theta_{i}\leftarrow\theta_{i}-\beta_{\theta}\nabla_{\theta_{i}}\mathcal{L}_{\text{SoftQ}}(\theta_{i})$
   if $\left(\text{iteration count}\mod d\right)=0$ then
@@
Algorithm 1 incorporates two implementation refinements. First, for the soft Q-loss (14), we replace $Q_{\theta}$ with $\min\{Q_{\theta_{1}},Q_{\theta_{2}}\}$ to mitigate overestimation bias. Second, we implement delayed updates for both policy $\pi_{\phi}$ and temperature $\alpha$, performing $d$ consecutive updates within each delayed cycle to ensure sufficient policy improvement. These techniques, adapted from the cleanrl framework [42], significantly improve training stability and performance.
@@ V Experiments
This section evaluates the MSACL algorithm across six benchmarking environments, covering stabilization and high-dimensional tracking tasks. We compare our approach against model-free and Lyapunov-based RL baselines using performance and stability metrics. Further analysis includes Lyapunov certificate visualization, robustness assessment under parametric uncertainty and stochastic noise, and sensitivity studies on the multi-step horizon $n$.
V-A Benchmarking Environments
Figure 3: Benchmarking environments. (a) VanderPol (Uncontrolled Phase Portrait), (b) Pendulum, (c) Planar DuctedFan, (d) Two-link Planar Robot, (e) SingleCarTracking, (f) QuadrotorTracking.
This subsection presents the benchmarking environments for evaluating MSACL and baseline algorithms. We categorize environments by control objective: stabilization tasks include Controlled VanderPol (VanderPol)[43], Pendulum[44], Planar DuctedFan (DuctedFan)[45], and Two-link planar robot (Two-link)[46]; tracking tasks comprise SingleCarTracking[47] and QuadrotorTracking[48], as illustrated in Figure 3.
For stabilization tasks, the goal is to drive the state to the origin $\mathbf{x}_{g}=\mathbf{0}\in\mathbb{R}^{d}$; for tracking tasks, the system follows a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$. All tasks use a unified reward structure:
@@ V-A Benchmarking Environments
where the first term penalizes state deviations and control effort using positive definite matrices $\mathbf{Q_{r}}$ and $\mathbf{R_{r}}$, minimizing the cumulative cost $\sum_{t=0}^{\infty}\left[\mathbf{x}_{t}^{\top}\mathbf{Q_{r}}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R_{r}}\mathbf{u}_{t}\right]$ used in optimal control [49]. For tracking tasks, $\mathbf{x}_{t}$ is replaced by the tracking error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$. The second term $r_{\text{approach}}$ provides sparse encouragement, defined as:
@@
Specifically, when $\|\mathbf{x}\|_{\infty}<\delta$ (e.g., 0.01), a positive reward $r_{\delta}$ is provided to incentivize convergence and mitigate steady-state error. All benchmarks follow the OpenAI Gym framework [50]. Notably, the Pendulum implementation is modified for increased difficulty, featuring expanded state and action spaces, boundary-based termination, and longer episodes, to better differentiate algorithmic performance. Specific details can be found in our code repository. During interaction, the environment returns a 5-tuple $(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$, with state transitions updated via explicit Euler integration.
@@ V-B Baseline Algorithms
We evaluate MSACL against classic model-free RL baselines (SAC [33], PPO [31]) and state-of-the-art Lyapunov-based algorithms (LAC [36], POLYC [25]). SAC and PPO serve as performance baselines for stabilization and tracking tasks, while LAC and POLYC provide stability guarantees through Lyapunov constraints. For LAC, the standard reward is replaced with a quadratic cost function defined as:
@@
where $\mathbf{Q_{c}}$ is positive definite. Note that in contrast to this state-only penalty formulation, our reward $r_{t}$ accounts for both state deviations and control effort. By comparing against these baselines, we aim to demonstrate the advantages of MSACL in terms of performance and stability metrics.
For fair comparison, all algorithms are implemented in the GOPS framework [51]. Actor and critic networks utilize MLPs with two 256-unit hidden layers and ReLU activations. All actors employ stochastic diagonal Gaussian policies. Critic architectures vary: LAC uses only a Lyapunov network, while POLYC incorporates both value and Lyapunov functions. SAC and LAC also implement clipped double Q-learning and delayed policy updates according to Remark 3. Optimization is performed via the Adam optimizer [30], with core hyperparameters listed in Table I.
@@ V-C Performance Evaluation and Stability Analysis
This subsection presents training outcomes, stability metrics, and Lyapunov certificate visualizations. Each algorithm is trained across five random seeds, with training results illustrated in Fig. 4 using the following performance metrics:
  • AMCR (Average Mean Cumulative Reward): The average of the Mean Cumulative Reward (MCR) computed across five independent training trials. MCR represents the total undiscounted reward accumulated per episode, normalized by the episode length to measure per-step performance.
  • AMCC (Average Mean Cumulative Cost): The average of the Mean Cumulative Cost (MCC) computed across five independent training trials. MCC represents the total undiscounted cost accumulated per episode, normalized by the episode length to quantify the average per-step magnitude of state deviations.
@@
Figure 4: Training performance across six benchmarks. The results show the average MCR and MCC over five independent runs. Solid lines and shaded regions represent the mean and one standard deviation, respectively. The horizontal axis indicates the number of training steps.
Training curves show MSACL outperforms or at least matches all baselines across six benchmarks. A positive AMCR indicates that the system state successfully enters the target $\delta$-neighborhood, while higher reward values imply faster convergence as the agent spends more time receiving $r_{\delta}$.
For comparison, we select the optimal policy for each algorithm, defined as the model that achieved the highest MCR among five training runs. These optimal policies are evaluated from 100 fixed initial states, with stability and convergence are assessed across four target radii ($0.2,0.1,0.05,0.01$) using the following stability metrics:
  • Reach Rate (RR): Percentage of trajectories that enter the specified radius.
  • Average Reach Step (ARS): Mean steps that required for the state to first enter the radius.
  • Average Hold Step (AHS): Mean steps that the state remains within the radius after the initial entry.
@@
TABLE II: Experimental Results of Baseline Algorithms Across Benchmarking Environments
@@
  • Note: AMCR and AMCC values represent the mean performance of the optimal policy (highest evaluation MCR) averaged over 100 independent trials, with $\pm$ denoting one standard deviation. Stability metrics (RR, ARS, and AHS) are evaluated across four radii; – indicates the radius was not reached. Bold entries highlight superior results: maximum AMCR, minimum AMCC, and highest RR. For stability metrics, the highest RR is prioritized for bolding, with the shortest ARS and longest AHS bolded for identical RR values.
Evaluation results demonstrate MSACL’s superiority. Specifically, MSACL consistently achieves the highest AMCR across all six benchmarks. Regarding AMCC, MSACL attains the absolute minimum in the Pendulum, SingleCarTracking, and QuadrotorTracking tasks, while remaining competitive in others. Furthermore, MSACL maintains the highest RR across all radii. Notably, in the high-dimensional QuadrotorTracking task, MSACL is the only approach to achieve a 100% reach rate at the $0.01$ precision level, whereas baselines fail. These results demonstrate the speed and precision of MSACL in complex control systems.
Finally, we visualize the Lyapunov networks $V_{\psi}$ for the optimal policies. The learned certificates and associated contour plots are shown in Fig. 5. For high-dimensional systems, the plots illustrate the relationship between $V_{\psi}$ and two specific state components with all other components set to zero.
@@
Since MSACL builds on the SAC framework, SAC serves as the primary baseline for our ablation study. As shown in Fig. 4 and Table II, the integration of the Lyapunov certificate significantly enhances both training performance and stability metrics compared to the standard SAC. This validates that embedding Lyapunov-based constraints into MERL effectively yields neural policies with superior stability guarantees for model-free control systems.
Figure 5: Visualization of learned Lyapunov certificates and their corresponding contour plots across six benchmarks: (a) VanderPol; (b) Pendulum; (c)–(e) DuctedFan (projections in the $(x,y)$-plane with $\theta=0$, $(y,\theta)$-plane with $x=0$, and $(x,\theta)$-plane with $y=0$); (f) Two-link; (g) SingleCarTracking; (h)–(j) QuadrotorTracking (projections in the $(e_{x},e_{y})$-plane with $e_{z}=0$, $(e_{y},e_{z})$-plane with $e_{x}=0$, and $(e_{x},e_{z})$-plane with $e_{y}=0$).
@@ V-D Evaluation of Robustness and Generalization
Neural policies with high-dimensional parameter spaces are prone to overfitting, often resulting in reduced robustness to noise and parametric variations, and limited generalization to unseen scenarios. This subsection evaluates the robustness and generalization of MSACL-trained policies, with the aim of demonstrating their real-world applicability under inherent uncertainties.
For stabilization tasks (VanderPol, Pendulum, DuctedFan, Two-link), we introduce process noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{env}}\cdot\mathbf{1}_{d})$ to the system dynamics, where $\mathbf{1}_{d}$ denotes the all-ones vector of dimension $d$, yielding the perturbed form:
@@
Additionally, we modify key physical parameters of the systems to assess whether the MSACL-trained policies can maintain stability under environmental variations. For tracking tasks (SingleCarTracking and QuadrotorTracking), robustness is evaluated via dynamical noise, while generalization is assessed using reference trajectories $\mathbf{x}_{t}^{\text{ref}}$ that are significantly different from those in training. Table III details the specific parameter modifications and noise magnitudes.
@@
Robustness results for stabilization tasks in Fig. 6 show that all trajectories converge to the equilibrium $\mathbf{x}_{g}=\mathbf{0}$ despite parametric perturbations and process noise. Tracking task results (Fig. 7, 8) demonstrate that MSACL-trained policies effectively mitigate process noise within reasonable ranges. For SingleCarTracking, though trained on straight-line references, the policy generalizes well to circular and sinusoidal trajectories (Fig. 7c, 7d), with position error $e_{\mathrm{xy}}=\sqrt{e_{x}^{2}+e_{y}^{2}}$ remaining below 0.1 in steady state (Fig. 7e). For QuadrotorTracking, though trained on horizontal helical trajectories, the policy successfully tracks unseen vertical helix and Lissajous curves (Fig. 8c, 8d), with tracking error $e_{\mathrm{xyz}}=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ stabilizing around 0.02 (Fig. 8e), demonstrating superior generalization capability.
Figure 6: System trajectories under the MSACL-trained policies in stabilization tasks subjected to environmental parametric uncertainties and process noise: (a) VanderPol with state variables $\mathbf{x}=[x,\dot{x}]^{\top}$; (b) Pendulum with angular position $\theta$ and angular velocity $\dot{\theta}$; (c) DuctedFan with coordinates $x,y$ and orientation $\theta$; (d) Two-link with joint angles $q_{1}$ and $q_{2}$. All trajectories exhibit robust convergence to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$.
@@
These results demonstrate that MSACL-trained policies exhibit robustness to model uncertainties and process noise, while generalizing effectively to unseen reference signals. However, performance degrades under extreme disturbances or significant deviations from the training distribution. In such cases, MSACL can be redeployed to retrain policies for stable, high-precision control in specific operational scenarios.
V-E Sensitivity Analysis of the Multi-Step Horizon
This subsection examines the impact of multi-step horizon $n$ on MSACL performance across all benchmarks. Following the protocol in Section V-C, we evaluate policies trained with varying $n$ using performance metrics (AMCR, AMCC) and stability metrics (RR, ARS, AHS). For each benchmark, we conduct 100 experiments from fixed random initial states to compute these metrics. As summarized in Table IV, performance generally improves with larger $n$, as a longer horizon enables the policy to better identify the stability direction, i.e., the descent of $V_{\psi}$. This reduces training bias and reinforces stability guarantees.
Empirically, $n=20$ delivers satisfactory performance across all benchmarks, serving as a versatile default. However, excessively large $n$ may degrade performance due to increased cumulative variance in training data, potentially distorting the optimization landscape as analyzed in Section III-B. Therefore, selecting an appropriate $n$ is essential to balance the bias-variance trade-off. For the control tasks herein, $n=20$ provides consistently robust results; for better performance, fine-tuning around this value is recommended.